Question - if BOTH Forces P1 and P2 were 10kips downward, then Member AB would be in COMPRESSION True False
TOTAL
Answer #1

. First calculate reaction RA+ R₂ = 20 kip [ {F4 = 0] EMA=0 .: Rix24-10X8-10X12 = 0 .. RLE 8.33 ↑ RA= 11-67 Now consider Join. First calculate reaction RA+ R₂ = 20 kip [ {F4 = 0] EMA=0 .: Rix24-10X8-10X12 = 0 .. RLE 8.33 ↑ RA= 11-67 Now consider Join